# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-edge AI Applications

### v1

@@ 6 Related Work
NVIDIA Collective Communications Library (NCCL) [21] and AMD RoCm Collective Communications Library (RCCL) [4] are vendor supplied libraries for NVIDIA and AMD GPUs respectively. Cowan et al. [9] improved over these libraries by allowing execution of collective algorithms specialized for a size and topology. In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
@@
The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication. For example, the NVIDIA OpenSHMEM (NVSHMEM) [22] is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective communication functions such as nvshmemx_broadcastmem_warp. However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and qualitative comparison is difficult as NVSHMEM is not open source. ARK [13] is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
@@
UCX, an open-source framework developed by Shamis et al. [33] and libraries implementing the MPI standard [35] offer a comprehensive set of APIs for high-performance computing (HPC) environments, however, MSCCL++ is more specific to GPU acceleration.
@@
Works like SCCL [6], TACCL [32], and TE-CCL [18] aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath. The primitives interface provided by MSCCL++ can be used with the generated algorithms to capture their benefit.
@@
Finally, LLM inference frameworks like vLLM [16] and TensorRT-LLM [25] implement custom AllReduce kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to single-node collective communication. MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
7 Conclusion
@@
MSCCL++ is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to 3.8×3.8$\times$3.8 × speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.

### v2

@@ 6. Related Work
NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025a) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively. Cowan et al. (Cowan et al., 2022) improved over these libraries by allowing execution of collective algorithms specialized for a size and topology. In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
@@
The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication. For example, the NVIDIA OpenSHMEM (NVSHMEM) (NVIDIA, 2025b) is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective functions such as nvshmemx_broadcastmem_warp. However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and quantitative comparison is difficult as NVSHMEM is not open source. ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
@@
UCX, an open-source framework developed by Shamis et al. (Shamis et al., 2015) and libraries implementing the MPI standard (Thakur et al., 2005) offer a comprehensive set of APIs for high-performance computing (HPC) environments. In contrast, MSCCL++ is more specific to GPU acceleration.
@@
SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath. The MSCCL++ primitives can be used with the generated algorithms to capture their benefit.
@@
Finally, LLM inference frameworks like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2025e) implement custom AllReduce kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to single-node collective communication. MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
7. Conclusion
@@
MSCCL++ is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to 5.4×5.4$\times$5.4 × speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.